# Exploring Token-Level Augmentation in Vision Transformer for Semi-Supervised Semantic Segmentation

### v1

@@ Abstract
Semi-supervised semantic segmentation has witnessed remarkable advancements in recent years. However, existing algorithms are based on convolutional neural networks and directly applying them to Vision Transformers poses certain limitations due to conceptual disparities. To this end, we propose TokenMix, a data augmentation technique specifically designed for semi-supervised semantic segmentation with Vision Transformers. TokenMix aligns well with the global attention mechanism by mixing images at the token level, enhancing learning capability for contexutual information among image patches. We further incorporate image augmentation and feature augmentation to promote the diversity of augmentation. Moreover, to enhance consistency regularization, we propose a dual-branch framework where each branch applies both image augmentation and feature augmentation to the input image. We conduct extensive experiments across multiple benchmark datasets, including Pascal VOC 2012, Cityscapes, and COCO. Results suggest that the proposed method outperforms state-of-the-art algorithms with notably observed accuracy improvement, especially under the circumstance of limited fine annotations.
@@ I Introduction
Recent research in S4 is mostly constructed upon Convolutional Neural Network (CNN)[15]. Although CNN has shown excellent performance in visual tasks, Vision Transformer (ViT)[16], as an emerging vision representation learning model, has demonstrated its superior performance and unique advantages in multiple computer vision tasks[5],[17],[18],[19]. Regarding the S4 task, directly applying algorithms to ViT specifically designed for CNN does not yield expected performance due to their conceptual differences[20]. Therefore, exploring how to effectively integrate ViT into the S4 framework to leverage its powerful global information processing capabilities requires in-depth research.
@@ II-B TokenMix
The structure of model $M$ can be divided into three parts: Patch Embedding, Encoder, and Decoder. Patch Embedding divides the image $x$ into patches and transforms them into high-dimensional tokens $g$. Encoder performs positional encoding on the tokens, and then extracts image features $f$ through self-attention and feedforward layers. Decoder calculates the final segmentation results $p$ of the image based on the features extracted by the Encoder. The process is shown in Fig. 2.
@@
where $m$ is a binary mask with 1 indicating that the corresponding parts should exchange tokens and 0 indicating that the corresponding parts should remain unchanged. After $g^{l}$ and $g^{u}$ are processed by Encoder, the tokens are then exchanged back:
@@
We do not use TokenMix on labeled images as this may affect the learning of annotation information, especially when the labels are scarce.
@@ III-B Comparisons With State-of-the-Arts
Figure 5: Effectiveness of TokenMix in larger models on Pascal.
@@ III-C Ablation Studies
Different Augmentations: We compare TokenMix with other augmentation methods in Table LABEL:augmentation, where our TokenMix achieves the best results. Combining TokenMix with other augmentations may reduce performance due to excessive transformations disrupting the original semantic structure. TokenMix in [34] (denoted as TokenMix*) is designed for fully supervised image classification, used for pre-training models that are fine-tuned for fully supervised semantic segmentation. Besides, TokenMix* swaps patches within several blocks at a fixed 0.5 ratio, while our TokenMix randomly swaps patches within a single block with a randomized ratio. Table LABEL:augmentation shows that our TokenMix significantly outperforms TokenMix* in S4.
Integration Into Other Methods: TokenMix can be seamlessly integrated into other SOTA methods to enhance their performance as shown in Table III.
@@ IV Conclusion
We reconsider data augmentation techniques for ViT-based S4 and propose TokenMix, which introduces new contextual information to help ViT better understand patch relationships. Leveraging ViT’s global processing strength, we further introduce a dual-branch framework that combines image and feature augmentation for greater diversity of augmentation. Experiments show our method is effective, particularly with limited labeled data.

### v2

@@ Abstract
Semi-supervised semantic segmentation has witnessed remarkable advancements in recent years. However, existing algorithms are based on convolutional neural networks and directly applying them to Vision Transformers poses certain limitations due to conceptual disparities. To this end, we propose TokenMix, a data augmentation technique specifically designed for semi-supervised semantic segmentation with Vision Transformers. TokenMix aligns well with the global attention mechanism by mixing images at the token level, enhancing learning capability for contextual information among image patches. We further incorporate image augmentation and feature augmentation to promote the diversity of augmentation. Moreover, to enhance consistency regularization, we propose a dual-branch framework where each branch applies image and feature augmentation to the input image. We conduct extensive experiments across multiple benchmark datasets, including Pascal VOC 2012, Cityscapes, and COCO. Results suggest that the proposed method outperforms state-of-the-art algorithms with notably observed accuracy improvement, especially under limited fine annotations.
@@ I Introduction
Recent research in S4 is mostly constructed upon Convolutional Neural Network (CNN)[15]. Although CNN has shown excellent performance in visual tasks, Vision Transformer (ViT)[16], as an emerging vision representation learning model, has demonstrated its superior performance and unique advantages in multiple computer vision tasks[5],[17],[18],[19]. However, algorithms developed for CNN work less effectively on ViT [20] due to their inherent architectural and conceptual disparities. Therefore, exploring how to effectively integrate ViT into S4 tasks to leverage its powerful global information processing capabilities requires in-depth research.
@@ II-B TokenMix
The structure of model $M$ can be divided into three parts: Patch Embedding, Encoder, and Decoder. Patch Embedding divides the image $x$ into patches and transforms them into high-dimensional tokens $g$. Encoder performs positional encoding on the tokens and then extracts image features $f$ through self-attention and feedforward layers. Decoder calculates the final segmentation results $p$ of the image based on the features extracted by the Encoder. The process is shown in Fig. 2.
@@
where $m$ is a binary mask where 1 indicates token exchange and 0 indicates no change. After $g^{l}$ and $g^{u}$ are processed by Encoder, tokens are then exchanged back:
@@ III-B Comparisons With State-of-the-Arts
Figure 5: Effectiveness of TokenMix in larger models.
@@ III-C Ablation Studies
Different Augmentations: We compare TokenMix with other augmentation methods in Table LABEL:augmentation, where our TokenMix achieves the best results. Combining TokenMix with other augmentations may reduce performance due to excessive transformations disrupting the original semantic structure. Note that [34] also presents token-level exchange (denoted as TokenMix*). However, it is designed for fully supervised image classification and used for pre-training models that are fine-tuned for fully supervised semantic segmentation, which our TokenMix is specifically designed for S4. Besides, Our TokenMix randomly swaps patches within a single block. In contrast, TokenMix* swaps patches within several small blocks, which introduces much inconsistent semantic information, making it difficult for the model to correctly model the true distribution of unlabeled data in S4. Table LABEL:augmentation shows that our TokenMix significantly outperforms TokenMix* in S4.
Integration Into Other Methods: TokenMix can be seamlessly integrated into other SOTA methods to enhance performance further, as shown in Table III.
@@ IV Conclusion
We reconsider data augmentation techniques for ViT-based S4 and propose TokenMix, which introduces new contextual information to help ViT better understand patch relationships. Leveraging ViT’s global processing strength, we further introduce a dual-branch framework that combines image and feature augmentation for greater diversity of augmentation. Experiments demonstrate our method’s effectiveness, particularly with limited labeled data.